# Strong constraints on a sharp change in $G$ as a solution to the Hubble tension

### v1

@@ 2.3 Geochronometry and cyclostratigraphy
The proposed drop in $G$ would increase the length of a year $\propto G^{2}$. Using the length of a day as a standard clock which would not change rapidly due to a drop in $G$, we see that the number of days per year would undergo a sharp increase on geological timescales. The minimum plausible 5% drop in $G$ would cause a 10% increase. This corresponds to each year having only 332 days prior to the transition and the modern 365 days afterwards.
@@
This unusual behaviour can be tested using geochronometry, the idea that fossilised remains of living organisms are sensitive to familiar cycles. It is common to use tree rings to reconstruct past climate conditions and count the age of a tree in years. Of more importance for our discussion is an underwater analogue to this idea. In particular, corals rely on sunlight, making them very sensitive to day/night cycles. But their growth also varies over the year due to seasons. By studying coral growth patterns in great detail, it is possible to determine the number of days per year many hundreds of Myr in the past (Wells, 1963; de Winter et al., 2020). The related technique of cyclostratigraphy exploits the related idea that deposition of sediments is also cyclic (Huang et al., 2024). [^1] Workers sometimes assume a fixed length of year and report the observed number of days per year as the number of hours per day.
These studies reveal that the number of days per year continuously declined with time. This is thought to be due to days gradually getting longer thanks to tidal evolution of the Earth–Moon system (Section 2.4). Tides would no doubt also operate in the GSM, but there would be an additional sharp 10% jump superimposed on the secular evolution. However, the geological data reveal only smooth trends, with the results from $>130$ Myr ago smoothly extrapolating onto the present value. The data are therefore in tension with a 10% jump in the number of days per year 130 Myr ago.
@@ 2.4 Impact on the Earth–Moon tidal evolution
The proposed drop in $G$ would enlarge not only Earth’s orbit around the Sun, but also the Moon’s orbital radius $R$ around the Earth. Following similar arguments, the GSM implies that $R\propto 1/G$. Since the lunar tidal stress on the Earth $\propto G/R^{3}\propto G^{4}$, a decrease in $G$ would substantially weaken oceanic tides on the Earth, which are largely caused by the Moon. On long timescales, gravitational attraction between the oceanic tidal bulges and the Moon causes it to slowly recede from Earth, a phenomenon which has been directly detected thanks to lunar laser ranging (Folkner et al., 2014). The orbital angular momentum of the Moon ultimately comes from the rotational angular momentum of the Earth, which correspondingly slows down and thus has longer days.
The lunar recession rate and the terrestrial spindown rate must have been much higher prior to the transition. A sharp change in the terrestrial spindown rate would lead to a discontinuity in the time derivative of the number of days per year, which would evolve much more rapidly prior to the transition. No such discontinuity is apparent in the data over the past few hundred Myr (de Winter et al., 2020; Huang et al., 2024). All this makes it unlikely that there was a significant discontinuity 130 Myr ago both in the number of days per year and in how quickly this was evolving with time (though the evolution is not completely uniform; see Huang et al., 2024).
@@ 3 Impact on stellar probes of the distance ladder
The distance ladder can be calibrated with many different types of star, whose luminosities exhibit a variety of dependences on the value of $G$. Thus, distance indicators calibrated to match each other within 40 Mpc would likely become discrepant beyond the proposed transition. For instance, Cepheid variable pulsations respond differently to main sequence stars (Jain et al., 2013; Sakstein, 2013), whose average luminosity is essentially what the surface brightness fluctuation (SBF) technique measures (Cantiello & Blakeslee, 2023). The TRGB magnitude (Anand et al., 2022, 2024) responds still differently (Sakstein et al., 2019), which by comparison with Cepheid distances has allowed workers to place $\approx 5-10\%$ constraints on $G/G_{0}$ (Desmond et al., 2019; Desmond & Sakstein, 2020). Geometric distances to megamasers would not be much affected (Pesce et al., 2020); as all but one of the megamasers considered by those authors are beyond the GSM transition, further constraints could be obtained by comparing megamaser distances to those obtained in ways that are sensitive to $G$.
@@
It has recently become possible to use the Fundamental Plane relation of elliptical galaxies (Djorgovski & Davis, 1987; Dressler et al., 1987) to constrain the local redshift gradient anchored to the distance to the Coma Cluster (Said et al., 2024). This lies well beyond the proposed transition, but we have to go across the transition to estimate the distance to Coma and thereby calibrate the zero-point of the Fundamental Plane distance-redshift relation. Different techniques give rather similar distances to Coma, with none suggesting that it lies $>110$ Mpc away (Scolnic et al., 2024). However, the local $cz^{\prime}=H_{0}^{\mathrm{Planck}}$ only if this is the case. For our discussion, the important point is the agreement between the distance to Coma as found using a variety of different distance indicators, whose absolute calibrations are generally performed closer than the proposed transition. The GSM would alter the SN distance to Coma in a manner that would solve the Hubble tension, but it seems very unlikely that the same can be said for all the other distance indicators. In particular, the luminosities of main sequence stars are much more sensitive to $G$ than the Tripp-corrected SN luminosity (Figure 1). Because of this, it seems difficult to place Coma $>110$ Mpc away while retaining agreement with the different available distance indicators, at least if the underestimated distances are due to underestimated $G$.
@@
Another issue with the GSM concerns the tight scaling relations evident in galaxies, especially the radial acceleration relation (RAR; Lelli et al., 2017) and the related Tully–Fisher (Tully & Fisher, 1977) relation in spiral galaxies and the Faber–Jackson (Faber & Jackson, 1976) relation and Fundamental Plane in elliptical galaxies. Regardless of the underlying origin of such relations, if the $M_{\star}/L$ ratios of galaxies within 40 Mpc are 30% higher than for more distant galaxies, then the scaling relations could not be as tight as they are observed to be. In addition, there would be substantial differences between samples split according to the transition distance. However, the intrinsic scatter in the RAR is very small (Li et al., 2018; Desmond, 2023) and the residuals do not correlate with a wide variety of galaxy properties, including distance (Stiskalek & Desmond, 2023). This is highly problematic for the GSM because the Spitzer Photometry and Accurate Rotation Curves (SPARC) dataset in which the RAR is best measured has many galaxies on either side of the proposed transition (Lelli et al., 2016).
@@ 3.1 Stellar ages and cosmic chronometers
The ages of the oldest stars and stellar populations in the Galactic disc and halo (Cimatti & Moresco, 2023) agree quite well with the age of the Universe in the Planck cosmology assumed by the GSM (e.g., Banik & Samaras, 2024). This agreement applies to standard calculations of the stellar ages. If instead the ages are recalculated assuming higher $G$ over nearly the entire lifetime of each star, then there would be an $\approx 3$ Gyr mismatch. It would be quite unusual if no stars were identifiable from the first 3 Gyr, especially when considering that several galaxies have already been discovered by the JWST at $z>14$ (Carniani et al., 2024). In the Planck cosmology assumed by the GSM, this corresponds to only 0.3 Gyr after the Big Bang, which may be faster than expected in the $\Lambda$CDM model (Haslbauer et al., 2022). A more severe version of this age gap problem applies to the covarying coupling constants model, which predicts that the universe is 26.7 Gyr old (Gupta, 2023).
It is also possible to constrain the expansion history using cosmic chronometers (CCs; Moresco et al., 2018, 2020). The basic idea is to get the time elapsed between two redshifts by considering ‘red and dead’ quiescent elliptical galaxies that presumably formed their stars at very early times. The stellar population of such a galaxy would evolve passively, with only the less massive stars remaining at later times. This would change the relative strengths of spectral lines, allowing a determination of the relative age between galaxies at the two considered redshifts. The time difference would be overestimated in the GSM because stars would evolve far faster at higher $G$, giving the appearance that much more time had elapsed. Observers assuming constant $G$ would then significantly underestimate the Hubble parameter, almost certainly making it no longer consistent with the predicted evolution in the assumed Planck background cosmology. However, the expansion history that it predicts is in excellent agreement with that reconstructed using the CC technique (Gómez-Valent & Amendola, 2018; Cogato et al., 2024). It is difficult to understand how this agreement can be preserved in the GSM, even if stellar luminosities scale with $G$ somewhat less steeply than expected theoretically ($G^{7}$; see Teller, 1948; degl’Innocenti et al., 1996). It is also very problematic that a change in $G$ would necessarily apply to both Hubble flow SNe and main sequence stars. This serious problem with the GSM could potentially be avoided in models with a screening mechanism, which we discuss in the next section.
@@ 4 Summary and model extensions
    The Solar luminosity $L_{\odot}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}G^{7}$ (Teller, 1948; degl’Innocenti et al., 1996). A sharp drop in $G$ and thus $L_{\odot}$ (combined with the resulting expansion of Earth’s orbit) would have plunged the Earth into a planetary glaciation due to the ice-albedo feedback effect (Budyko, 1969; Sellers, 1973). Although there have been several Snowball Earth episodes, none occurred in the time period relevant to the GSM (Banik, 2016, and references therein).
@@
    A reduction in $G$ would cause Earth’s orbit to expand, increasing the length of a year. However, the number of days would not change in like manner because the Earth is bound by chemical forces rather than gravity. As a result, the number of days per year would sharply increase by $\ga 10\%$. It is possible to constrain this using geochronometry and cyclostratigraphy (Section 2.3), for instance using corals that are sensitive to both daily and seasonal cycles (Wells, 1963). There is no indication of such a rapid change in the number of days per year (de Winter et al., 2020; Huang et al., 2024). Instead, extrapolating trends in the data prior to the $G$ transition gives a value close to the present-day value of 365.25.
@@
    The expansion of gravitationally bound orbits would also apply to the Moon. A rapid increase in the Moon’s distance from Earth combined with a reduced $G$ would sharply decrease the amplitude of Earth’s oceanic tidal bulges. Since gravitational interactions between these bulges and the Moon cause the Earth’s spin to slow down, there would be a discontinuity in the time derivative of the number of days per year, which must have evolved much more rapidly prior to the transition. There is again no evidence of this, suggesting that the distance to the Moon did not jump by 10% on a geologically short timescale in the last few hundred Myr.
@@
    The local $cz^{\prime}$ can be obtained using several other techniques besides the traditional Cepheid–SNe route (e.g. Riess & Breuval, 2024, and references therein). Since these give quite similar results, it is unlikely that the evolution of $G$ can be chosen to alter all the distances in step with each other. In particular, the GSM implies that stars were substantially brighter prior to the proposed transition. This would lead to a very severe mismatch between distance techniques that rely on stellar luminosities and those that do not. For instance, we would expect a severe mismatch between redshift distances and SBF distances, which are sensitive to the luminosity-weighted mean stellar luminosity (Cantiello & Blakeslee, 2023). A drastic difference between the $M_{\star}/L$ ratios of galaxies within and beyond the proposed transition would also introduce a large mismatch in galaxy scaling relations like the Tully–Fisher relation and Fundamental Plane, in particular making it very difficult to understand how galaxy samples that extend both sides of any plausible transition radius follow such a tight RAR (Desmond, 2023; Stiskalek & Desmond, 2023).
@@
    The GSM implies substantially faster stellar evolution prior to the transition, and thus over the vast majority of cosmic history. This would lead to the appearance that stellar populations of quiescent galaxies have aged much more between any two redshifts, causing the cosmic expansion history estimated using the CC technique to differ substantially from the Planck cosmology assumed in the GSM. However, there is excellent agreement in practice (Cogato et al., 2024). This also applies to the ages of the oldest stars and stellar populations in the Galactic disc and halo, which yield a cosmic age in good agreement with the Planck cosmology (Cimatti & Moresco, 2023). The GSM would force a substantial downward revision to these ages, leading to an age gap of $\approx 3$ Gyr in the early universe lacking any known stellar counterpart.
@@
Our consideration of these issues suggests that the GSM faces insurmountable difficulties when additional constraints are imposed on the evolution of $G$ over the past $\sim$300 Myr, the light travel time to the nearest Hubble flow SNe. Moreover, the local redshift gradient can be obtained in several ways that do not rely on the traditional Cepheid–SNe route, yet give similar results. We therefore argue that the GSM is not a viable solution to the Hubble tension.
@@
The GSM focuses on the possibility that the local $cz^{\prime}$ has been overestimated because of a mismatch between the luminosities of SNe in the Hubble flow and those in more nearby host galaxies with an independent Cepheid or TRGB calibration. The difficulties with the GSM lead us to consider if the problem might lie rather with the first rung, which assumes that the Leavitt law and TRGB magnitude are the same between anchor galaxies and SN host galaxies. One way to violate this assumption would be through a rapid evolution in $G$, which would need to occur even more recently than in the GSM. This would if anything exacerbate the already very severe problems we have identified with the GSM.

### v2

@@ 2.3 Geochronometry and cyclostratigraphy
The proposed drop in $G$ would increase the length of a year (LOY) $\propto G^{-2}$. However, the length of a day (LOD) would barely change because the Earth’s radius $R_{\oplus}$ would not increase much. Since angular momentum conservation applies both to Earth’s orbit around the Sun and to its daily rotation on its own axis, we would require $R_{\oplus}\propto G^{-1}$ to preserve the ratio LOY/LOD. Since volume scales as radius cubed, a 5% drop in $G$ would have to cause a 15% drop in Earth’s average density $\rho_{\oplus}$. This is unrealistic because gravitational compression has increased the uncompressed density of the Earth by only 36%, from 4050 to 5515 kg/m3 (Hughes, 2006). Thus, reducing $G$ by a mere 5% cannot be expected to cause a 15% drop in $\rho_{\oplus}$, especially given that Earth’s interior structure and temperature were quite similar to present values 130 Myr ago as this represents only a small fraction of its age (Herzberg et al., 2010). Another way to see this is to note that Earth rotates much slower than the Keplerian orbital velocity at its surface, so the vast majority of the support against gravity is provided by pressure gradients in nearly incompressible molten rock. As a result, reducing $G$ cannot cause Earth to expand as much as would a purely rotationally supported system. Since gravity is responsible for increasing $\rho_{\oplus}$ by 36%, a more reasonable assumption is that a 5% drop in $G$ would reduce $\rho_{\oplus}$ by 1.8% and thus increase $R_{\oplus}$ by only 0.6%, almost an order of magnitude smaller than the required 5%. Thus, we can use the LOD as a standard clock which would not change rapidly due to a drop in $G$, unlike the LOY. This implies that the number of days per year would undergo a sharp increase on a geologically short timescale. The minimum plausible 5% drop in $G$ would cause a 10% increase. This corresponds to each year having only 332 days prior to the transition and the modern 365 days afterwards.
@@
This unusual predicted behaviour can be tested using geochronometry, the idea that fossilised remains of living organisms are sensitive to familiar cycles. It is common to use tree rings to reconstruct past climate conditions and count the age of a tree in years. Of more importance for our discussion is an underwater analogue to this idea. In particular, corals rely on sunlight, making them very sensitive to day/night cycles. But their growth also varies over the year due to seasons. By studying coral growth patterns in great detail, it is possible to determine the number of days per year many hundreds of Myr in the past (Wells, 1963; de Winter et al., 2020). The related technique of cyclostratigraphy exploits the related idea that deposition of sediments is also cyclic (Huang et al., 2024). Those authors report their results in terms of the LOD assuming a fixed LOY, but the actual observable is the ratio LOY/LOD, so the predicted rise in LOY would appear as a drop in their reported LOD.
Figure 2: Evolution of the LOD from geochronometry and cyclostratigraphy as a function of lookback time, assuming a fixed LOY. The results shown here are thus inversely proportional to the observable quantity LOY/LOD, which is found from counting the number of daily cycles per annual cycle. In the GSM, the apparent LOD would have abruptly decreased by $\ga 10\%$ about 130 Myr ago (dashed red line), as indicated by the downward red arrow. The evolution at other times is due to lunar tides, as captured to some extent by the models (solid lines). Geochronometry from the last 75 Myr (de Winter et al., 2020) and lunar laser ranging indicate that the LOD is presently increasing by 1 h every 150 Myr, in line with the evolution prior to the proposed transition. Reproduced from figure 4b of Huang et al. (2024).
For ease of reference, we reproduce figure 4b of Huang et al. (2024) in our Figure 2. In line with other studies, this reveals that the number of days per year continuously declined with time. This is thought to be due to days gradually getting longer thanks to tidal evolution of the Earth–Moon system (Section 2.4). Tides would no doubt also operate in the GSM, but there would be an additional sharp 10% jump superimposed on the secular evolution. However, the geological data reveal only smooth trends, with the results from $>130$ Myr ago smoothly extrapolating onto the present value. The data are therefore in tension with a 10% jump in the number of days per year 130 Myr ago, which would appear on Figure 2 as a 10% drop in the apparent LOD if assuming a fixed LOY. Even if we assume fairly rapid tidal evolution within the gap in the data $100-200$ Myr ago, it is clear that the apparent LOD would still have been $<24$ h 130 Myr ago, at which point it would drop 10% to $<22$ h, as indicated by the downward red arrow. Unusually rapid tidal evolution would be required at later times to achieve agreement with our present 24 h days. This is problematic because the present rate of tidal evolution is constrained from lunar laser ranging, which is already used as a constraint on the geological models shown in Figure 2. Given the similar positions of continents and oceans in the last 100 Myr, there is no good reason for the average rate of tidal dissipation over this period to differ much from its present rate, which if extrapolated backwards implies $>23$ h days 130 Myr ago. Moreover, as we discuss next, the larger Earth-Moon distance after the transition and consequent reduced tidal stress on Earth would if anything cause tidal evolution to be slower rather than faster after the proposed drop in $G$, leaving the apparent LOD stuck at $\la 22$ h and thus making each year $\ga 400$ days. Therefore, although the gap in the geological record near the vertical dashed red line in Figure 2 prevents us from drawing stronger conclusions about whether a 10% discontinuity in LOY/LOD is feasible 130 Myr ago, it does seem very unlikely that tidal evolution could reconcile such a discontinuity with available geological data.
@@ 2.4 Impact on the Earth–Moon tidal evolution
The proposed drop in $G$ would enlarge not only Earth’s orbit around the Sun, but also the Moon’s orbital radius $R$ around the Earth. Following similar arguments, the GSM implies that $R\propto 1/G$. Since the lunar tidal stress on the Earth $\propto G/R^{3}\propto G^{4}$, a decrease in $G$ would substantially weaken oceanic tides on the Earth, which are largely caused by the Moon. On long timescales, gravitational attraction between the oceanic tidal bulges and the Moon causes it to slowly recede from Earth, a phenomenon which has been directly detected thanks to lunar laser ranging (Folkner et al., 2014). The Moon pulls on both tidal bulges, but the bulge near it is more important due to the inverse square dependence of the gravity law. This bulge is dragged ‘ahead’ of the Moon because Earth rotates on its axis much faster than the Moon orbits it, leading to a net tidal torque. The orbital angular momentum of the Moon ultimately comes from the rotational angular momentum of the Earth, which correspondingly slows down and thus has longer days.
The lunar recession rate and the terrestrial spindown rate must have been much higher prior to the transition. A sharp change in the terrestrial spindown rate would lead to a discontinuity in the time derivative of the number of days per year, which would evolve much more rapidly prior to the transition. However, Figure 2 shows that the spindown rate at that time (left of red line) was about the same as it is today based on lunar laser ranging, which indicates that the LOD is increasing by 2.4 ms per century or 1 h every 150 Myr (Williams & Boggs, 2016). It is also possible to get the present spindown rate from geochronology, which indicates that the LOD was about 0.5 h shorter than today 75 Myr ago (de Winter et al., 2020). Therefore, a discontinuity in the spindown rate 130 Myr ago is not apparent in the data over the past few hundred Myr (de Winter et al., 2020; Huang et al., 2024). All this makes it unlikely that there was a significant discontinuity 130 Myr ago both in the number of days per year and in how quickly this was evolving with time (though the evolution is not completely uniform; see Huang et al., 2024).
@@ 3 Impact on stellar probes of the distance ladder
It has recently become possible to use the Fundamental Plane relation of elliptical galaxies (Djorgovski & Davis, 1987; Dressler et al., 1987) to constrain the local redshift gradient anchored to the distance to the Coma Cluster (Said et al., 2024). This lies well beyond the proposed transition, but we have to go across the transition to estimate the distance to Coma and thereby calibrate the zero-point of the Fundamental Plane distance-redshift relation. Different techniques give rather similar distances to Coma, with none suggesting that it lies $>110$ Mpc away (Scolnic et al., 2025). However, the local $cz^{\prime}=H_{0}^{\mathrm{Planck}}$ only if this is the case. For our discussion, the important point is the agreement between the distance to Coma as found using a variety of different distance indicators, whose absolute calibrations are generally performed closer than the proposed transition. The GSM would alter the SN distance to Coma in a manner that would solve the Hubble tension, but it seems very unlikely that the same can be said for all the other distance indicators. In particular, the luminosities of main sequence stars are much more sensitive to $G$ than the Tripp-corrected SN luminosity (Figure 1). Because of this, it seems difficult to place Coma $>110$ Mpc away while retaining agreement between the different available distance indicators, at least if the underestimated distances are due to underestimated $G$.
@@
Another issue with the GSM concerns the tight scaling relations evident in galaxies, especially the radial acceleration relation (RAR; Lelli et al., 2017) and the related Tully–Fisher (Tully & Fisher, 1977) relation in spiral galaxies and the Faber–Jackson (Faber & Jackson, 1976) relation and Fundamental Plane in elliptical galaxies. Regardless of the underlying origin of such relations, if the $M_{\star}/L$ ratios of galaxies within 40 Mpc are 30% higher than for more distant galaxies, then these scaling relations could not be as tight as they are observed to be. In particular, there would be substantial differences between samples split according to the transition distance. However, the intrinsic scatter in the RAR is very small (Li et al., 2018; Desmond, 2023) and the residuals do not correlate with a wide variety of galaxy properties, including distance (Stiskalek & Desmond, 2023). While their analysis did not specifically investigate a transition at a given distance, the machine learning techniques it utilised would be capable of picking out any significant variation. This is highly problematic for the GSM because the Spitzer Photometry and Accurate Rotation Curves (SPARC) dataset in which the RAR is best measured has many galaxies on either side of the proposed transition (see figure 1 of Lelli et al., 2016). Moreover, the hint for a transition in $G$ in the Tully–Fisher relation identified by Alestas, Antoniou & Perivolaropoulos (2021) is in the opposite direction to that required by the GSM. This was not mentioned in that study due to its neglect of the effect of standardisation on the $L-G$ relation of SNe (Wright & Li, 2018).
@@ 3.1 Stellar ages and cosmic chronometers
The ages of the oldest stars and stellar populations in the Galactic disc and halo (Cimatti & Moresco, 2023) agree quite well with the age of the Universe in the Planck cosmology assumed by the GSM (e.g., Banik & Samaras, 2024; Mazurenko et al., 2025). This agreement applies to standard calculations of the stellar ages. If instead the ages are recalculated assuming higher $G$ over nearly the entire lifetime of each star, then there would be an $\approx 3$ Gyr mismatch (degl’Innocenti et al., 1996). It would be quite unusual if no stars were identifiable from the first 3 Gyr, especially when considering that several galaxies have already been discovered by the JWST at $z>14$ (Carniani et al., 2024). In the Planck cosmology assumed by the GSM, this corresponds to only 0.3 Gyr after the Big Bang, which may be faster than expected in the $\Lambda$CDM model (Haslbauer et al., 2022). A more severe version of this age gap problem applies to the covarying coupling constants model, which predicts that the universe is 26.7 Gyr old (Gupta, 2023).
It is also possible to constrain the expansion history using cosmic chronometers (CCs; Moresco et al., 2018, 2020). The basic idea is to get the time elapsed between two redshifts by considering ‘red and dead’ quiescent elliptical galaxies that presumably formed their stars at very early times. The stellar population of such a galaxy would evolve passively, with only the less massive stars remaining at later times. This would change the relative strengths of spectral lines, allowing a determination of the relative age between galaxies at the two considered redshifts. Of particular importance to the CC technique is the D4000 break at 400 nm (Moresco, 2024). Calculations would need to be conducted of how this feature evolves with time at higher $G$. More generally, we would need to obtain a revised relation between the colour of a star and its lifetime. Although higher $G$ reduces the lifetime at fixed mass, it is possible that the colour also changes in such a way as to largely move stars along the standard colour-lifetime relation. Even so, it seems unlikely that such cancellation of different effects would be precise enough to explain the excellent agreement between the Planck background cosmology and that reconstructed using the CC technique (Gómez-Valent & Amendola, 2018; Cogato et al., 2024; Guo et al., 2024). The latter authors found that $H_{0}=H_{0}^{\mathrm{Planck}}$ to within 1 km/s/Mpc, which is also their stated uncertainty. It is difficult to understand how this agreement can be preserved in the GSM, even if stellar luminosities scale with $G$ somewhat less steeply than expected theoretically ($G^{5.6}$; see degl’Innocenti et al., 1996). It therefore seems very problematic that a change in $G$ would necessarily apply to both Hubble flow SNe and main sequence stars. This serious problem with the GSM could potentially be avoided in models with a screening mechanism, which we discuss in the next section.
@@ 4 Summary and model extensions
    The Solar luminosity $L_{\odot}\mathrel{\vbox{ \offinterlineskip\halign{\hfil$#$\cr\propto\cr\kern 2.0pt\cr\sim\cr\kern-2.0pt\cr}}}G^{5.6}$ (degl’Innocenti et al., 1996). If $G$ were higher for most of the Sun’s history, it would have consumed much more of its fuel in the precisely known time since it formed (Connelly et al., 2012). This would inflate its helioseismic age to well over $5.5$ Gyr, which is not the case observationally (Bétrisey et al., 2024).
@@
    A sharp drop in $G$ and thus $L_{\odot}$ (combined with the resulting expansion of Earth’s orbit) would likely have plunged the Earth into a planetary glaciation due to the ice-albedo feedback effect (Budyko, 1969; Sellers, 1973). Although there have been several Snowball Earth episodes, none occurred in the time period relevant to the GSM (Banik, 2016, and references therein).
@@
    A reduction in $G$ would cause Earth’s orbit to expand, increasing the LOY. However, the LOD would not change in like manner because the Earth is bound by chemical forces rather than gravity. Compression of the Earth by gravity only raises its density by 36% (Hughes, 2006), so a reduction in $G$ by 5% cannot be expected to reduce its overall density by 15%, which however would be needed to increase its radius by 5% in order to increase the LOD in line with the LOY. As a result, the number of days per year (LOY/LOD) would sharply increase by $\ga 10\%$. It is possible to constrain this using geochronometry and cyclostratigraphy (Section 2.3), for instance using corals that are sensitive to both daily and seasonal cycles (Wells, 1963). There is no indication of such a rapid change in the number of days per year (de Winter et al., 2020; Huang et al., 2024). Instead, extrapolating trends in the data prior to the $G$ transition gives a value close to the present-day value of 365.25. A gap in the geological data $100-200$ Myr ago prevents a clearer assessment of whether the sharp change predicted by the GSM actually occurred, though quite unusual tidal evolution of the Earth-Moon system would be required in this period to mask the effects predicted by the GSM (Figure 2).
@@
    The expansion of gravitationally bound orbits would also apply to the Moon. A rapid increase in the Moon’s distance from Earth combined with a reduced $G$ would sharply decrease the amplitude of Earth’s oceanic tidal bulges. Since gravitational interactions between these bulges and the Moon cause the Earth’s spin to slow down, there would be a discontinuity in the time derivative of the number of days per year, which must have evolved much more rapidly prior to the transition. There is again no evidence of this, with lunar laser ranging and geochronometry within the last 100 Myr indicating that the LOD is increasing by about 1 h every 150 Myr, in line with the evolution prior to the transition (Figure 2). This suggests that the distance to the Moon did not jump by 10% on a geologically short timescale in the last few hundred Myr.
@@
    The local $cz^{\prime}$ can be obtained using several other techniques besides the traditional Cepheid–SNe route (e.g. Riess & Breuval, 2024, and references therein). Since these give quite similar results, it is unlikely that the evolution of $G$ can be chosen to alter all the distances in step with each other. In particular, the GSM implies that stars were substantially brighter prior to the proposed transition. This would lead to a very severe mismatch between distance techniques that rely on stellar luminosities and those that do not. For instance, we would expect a severe mismatch between redshift and surface brightness fluctuation (SBF) distances, which are sensitive to the luminosity-weighted mean stellar luminosity (Cantiello & Blakeslee, 2023). A drastic difference between the $M_{\star}/L$ ratios of galaxies within and beyond the proposed transition would also introduce a large mismatch in galaxy scaling relations like the Tully–Fisher relation and Fundamental Plane. It is particularly difficult to understand how galaxy samples that extend both sides of any plausible transition radius follow such a tight RAR (Desmond, 2023), especially given that a model-agnostic machine learning-based approach to checking if residuals from the RAR correlate with heliocentric distance found no discernible correlation (Stiskalek & Desmond, 2023).
  6. 6.
@@
    The GSM implies substantially faster stellar evolution prior to the transition, and thus over the vast majority of cosmic history. This would cause the stellar populations of quiescent galaxies to evolve much more between any two redshifts, which would have significant effects on reconstructions of the expansion history using CCs. Higher $G$ would however also affect other aspects of a star at any fixed mass that might be used to estimate its age, for instance its colour and the strength of the D4000 feature in its spectrum. Even so, it is unlikely that these effects would balance so precisely that the CC technique would recover $H_{0}=H_{0}^{\mathrm{Planck}}$ to within 1 km/s/Mpc, as occurs in practice (Cogato et al., 2024; Guo et al., 2024). Moreover, the proposed higher $G$ in the past would not apply to current observations of the oldest stars and stellar populations in the Galactic disc and halo given their proximity. Estimating their ages assuming constant $G$ yields a cosmic age in good agreement with the Planck cosmology (Cimatti & Moresco, 2023; Banik & Samaras, 2024; Mazurenko et al., 2025). The GSM would force a substantial downward revision to these ages because these stars would have experienced higher $G$ for most of their history, leading to an age gap of $\approx 3$ Gyr in the early universe lacking any known stellar counterpart.
@@
Our consideration of these issues suggests that the GSM faces severe difficulties when additional constraints are imposed on the evolution of $G$ over the past $\approx 300$ Myr, the light travel time to the nearest Hubble flow SNe. Moreover, the local redshift gradient can be obtained in several ways that do not rely on the traditional Cepheid–SNe route, yet give similar results. We therefore argue that the GSM is not a viable solution to the Hubble tension.
@@
The GSM focuses on the possibility that the local $cz^{\prime}$ has been overestimated because of a mismatch between the luminosities of SNe in the Hubble flow and those in more nearby host galaxies with an independent Cepheid or TRGB calibration. The difficulties with the GSM lead us to consider if the problem might lie rather with the first rung, which assumes that the Leavitt law and TRGB magnitude are the same between geometric anchor galaxies and SN host galaxies. One way to violate this assumption would be through a rapid evolution in $G$, which would need to occur even more recently than in the GSM. This would if anything exacerbate the already very severe problems we have identified with the GSM. On the other hand, comparison of Cepheid and SNe distances within 30 Mpc provides some hints for a change in $G$ within the last 100 Myr based on a distance-dependent transition in the SN Ia absolute magnitude, exploiting the different way in which $G$ affects SNe and Cepheid variables (Perivolaropoulos & Skara, 2021; Ruchika et al., 2024a). In particular, Ruchika et al. (2024b) found evidence for a model in which $G$ decreased by 4% 80 Myr ago. However, their analysis requires too strong a $G$ dependence of the standardised SN luminosity. Those authors parametrise the dependence in the form $L\propto G^{-3n/2}$, with the theoretical expectation being $n=-0.97$ (Wright & Li, 2018). However, the joint inference on $n$ and $H_{0}$ is inconsistent with $n=-0.97$ and $cz^{\prime}=H_{0}^{\mathrm{Planck}}$ at well over $2\sigma$ confidence (see their figure 5). Indeed, their most preferred model has $n=-4.24\pm 1.06$, which is $>3\sigma$ discrepant with the theoretical expectation. Some of their models can largely resolve the Hubble tension if we allow $n\approx-2$, but this would require significant errors in the calculations of Wright & Li (2018). Moreover, a sharp change in $G$ at the reduced distance of 25 Mpc would exacerbate the problem discussed earlier concerning the tightness of the RAR and lack of residual correlations with heliocentric distance (Stiskalek & Desmond, 2023), as 25 Mpc is closer than 40 Mpc to the median distance of the SPARC sample.